Circular No. 7856

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1908+094 P. M. Woods, C. Kouveliotou, M. H. Finger, and E. Gogus, National

Space Science and Technology Center; and J. Swank, C. Markwardt, and T. Strohmayer, Goddard Space Flight Center, report the serendipitous discovery of a new x-ray transient, XTE J1908+094, in RXTE PCA observations of the soft-γ-ray repeater SGR 1900+14, triggered following the burst activity on Feb. 17–18 (GCN 1253). These observations failed to detect the 5.2-s SGR pulsations, pointing towards a possible new source as the origin of the high x-ray flux. An RXTE PCA scan of the region around SGR 1900+14 on Feb. 21 was consistent with emission only from known sources (and no new sources). However, the scans required SGR 1900+14 to be 20 times brighter than its quiescent flux level (GCN 1256). A Director’s Discretionary Time Chandra observation on Mar. 11 showed that the SGR was quiescent and did not reveal any new source within the Chandra ACIS field-of-view. A subsequent RXTE PCA scan on Mar. 17, taken in combination with the first scan, required that a new source be included in the fit. The best-fit position is α = 19^{h}08^{m}50^{s}, δ = +9^{o}22′.5 (equinox J2000.0; estimated 2′ systematic error radius), or ≈ 24′ away from the SGR source. The source spectrum (2–30 keV) can be best fit with a power-law function including photoelectric absorption (column density N_{h} = 2.3 × 10^{22}, photon index = 1.55). Iron line emission is present, but may be due to the Galactic ridge. Between Feb. 19 and Mar. 17, the source flux (2–10 keV) has risen from 26 to 64 mCrab. The power spectrum is flat between 1 mHz and 0.1 Hz, falling approximately as 1/f^{0.5} up to 1 Hz. At 1 Hz is seen a broad quasiperiodic oscillation peak and a break to a 1/f^{2} power law, which continues to 4 Hz. The fractional rms amplitude from 1 mHz to 4 Hz is 43 percent. No coherent pulsations are seen between 0.001 and 1024 Hz. The authors conclude that XTE J1908+094 is a new blackhole candidate.

γ-RAY TRANSIENT M. Schmidt, California Institute of Technology, writes: “The two γ-

ray transients observed by Konus and Ulysses on 1995 Jan. 10 and Mar. 25 (IAUC 7840) are clearly present in BATSE DISCLA data from GRO Compton. In a systematic search for γ-ray bursts in DISCLA data completed in 2001, it was found that three software triggers occurred during the Jan. 1995 event and two during the Mar. event. The localizations of the triggers are all within 17^{o} of Cyg X-1, consistent with the conclusion on IAUC 7840 that these two outbursts originated from Cyg X-1.”

2002 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green